OpenStudy (anonymous):

FILL IN THE BLANK 8 x (9+4)= ( x 9) + ( 8 x ) (4 x 7)+ ( 4 x 5)= 4 ( + 5) 9 x ( 7 -1 ) = ( 9 x ) -( x 1) help me solve the blanks

OpenStudy (anonymous):

i will give a medal

OpenStudy (ipwnbunnies):

This an exercise of the Distributive Property of Multiplication. It's like this: A*(B + C) = A*B + A*C

OpenStudy (ipwnbunnies):

Where A, B, and C are numbers. It also works for subtraction. So for the first question: 8*(9 + 4) = 8*9 + 8*4